# A Lightweight Transformer for Pain Recognition from Brain Activity

### 13 Apr 2026 v1

@@ IV-B Representation Fusion
The effect of combining complementary fNIRS representations is examined by stacking the waveform and PSD tensors along the channel dimension, thereby enabling joint modeling with a single shared backbone. For this analysis, $S=32$ segments are selected, as this configuration provides strong unimodal performance with balanced computational and inference costs. The fusion results are reported in Table IV. Stacking waveform and PSD achieves $48.52\%$ accuracy, $45.62\%$ precision, and $44.62\%$ F1 score, indicating that temporal dynamics in the raw signal and time–frequency structure in the PSD encode complementary information relevant to pain perception. The increase in computational cost remains modest, comparable to PSD-only processing, while inference latency remains within the same efficient regime on both GPU and CPU. To enable representation fusion, the 1D waveform grid is projected onto a 2D layout through interpolation and concatenated with the PSD tensor along the channel dimension:
@@ V Comparison with Existing Methods
Table V presents a comparison of the AI4Pain testing set across video-based, fNIRS-based, and multimodal approaches. Classical handcrafted-feature pipelines have limited capacity to capture complex pain-related patterns, with the SVM framework in [ai4pain_2024] achieving at most $43.30\%$ accuracy across modalities. Deep learning methods yield consistent improvements: the multimodal transformer in [gkikas_tsiknakis_painvit_2024] achieves $46.67\%$, followed by a video-only CNN at $49.00\%$ [prajod_schiller_2024], and a CNN–Transformer fusion at $51.33\%$ [vianto_2025]. The ensemble classifier in [khan_aziz_2025] reaches $53.66\%$ using handcrafted fNIRS features, while the video transformer in [nguyen_yang_2024] achieves $55.00\%$. The multimodal transformer in [gkikas_rojas_painformer_2025] reports the strongest result at $55.69\%$ through joint video–fNIRS learning. The proposed method achieves $50.00\%$ accuracy on the testing set using stacked waveform and PSD fNIRS representations. Relying exclusively on cortical signals without any behavioral input, this result surpasses all classical fNIRS pipelines and remains competitive with multimodal systems, supporting the viability of unified tokenization and representation fusion for fNIRS-based pain recognition.
@@ VI Conclusion
This work presented a lightweight transformer for fNIRS-based pain recognition, built around a unified tokenization and latent segmentation strategy that supports multi-representation fusion without architectural specialization. Evaluated on the AI4Pain dataset, the model achieves stable recognition performance across a broad range of computational configurations, with segment numbers between $2$ and $16$ offering the most favorable balance between efficiency and accuracy. Fusing raw waveform and PSD representations shows strong performance and confirms that temporal and time–frequency structures capture complementary information relevant to cortical pain dynamics. However, further exploration is needed to better understand the strengths and limitations of different representations. Despite relying only on brain activity and operating under tight computational constraints, the proposed method surpasses classical fNIRS pipelines and remains competitive with multimodal deep learning systems that incorporate behavioral signals. These results support the effectiveness of efficient transformer designs for neurophysiological pain assessment in constrained settings [gkikas_tiny_2025, antonogiorgakis_britzolakis_2019]. This is relevant to clinical and interactive environments [chatziadam_dimitraidis_2020, 10807220], where social robots could use real-time estimates of pain and cognitive states to adapt their interactions with users [arzate_when_2025, arzate_hri_2025, arzate_3dmm_2022, fang_moleron_2024, kruger_oshima_2026, hessels_fang_2026].

### 23 Apr 2026 v2

@@ IV-B Representation Fusion
The effect of combining the two fNIRS representations is evaluated by stacking the waveform and PSD tensors along the channel dimension. This enables a joint modeling with a single shared backbone. For this analysis, $S=32$ segments are selected, as this configuration provides strong unimodal performance with balanced computational and inference costs. The fusion results are reported in Table IV.
Using the stacking approach for the waveform and PSD representations led to $48.52\%$ accuracy, $45.62\%$ precision, and $44.62\%$ F1 score. Although this results in performance better than the highest performing PSD-only model ($52.31\%$ with $S=2$), we demonstrate that both representations may be used together as part of a common backbone for modeling without changing the underlying architecture. We further show that by combining representations using our method, the raw waveform does not provide additional information about the signal beyond what is contained in its PSD.
We observe that the increase in computational cost resulting from adding the waveform representation remains small compared to PSD-only processing, and similarly, inference latency has remained very similar to the PSD-only case on both the GPU and CPU. The representation fusion is perfomed by using the 1D waveform grid which is projected onto a 2D layout through interpolation and concatenated with the PSD tensor along the channel dimension:
@@ V Comparison with Existing Methods
In Table V we compare the performance of all three approaches using data from the AI4Pain test set: Video-based, fNIRS-based, and multimodal approaches. Handcrafted feature-based traditional methods were shown to have limited capacity for representing complex patterns associated with pain; even so, the best accuracy reported by an SVM approach presented in [ai4pain_2024] was $43.30\%$, regardless of modality. From the othe hand deep learning methods generally perform better. The multimodal transformer in [gkikas_tsiknakis_painvit_2024] achieves $46.67\%$, the video-only CNN in [prajod_schiller_2024] reaches $49.00\%$, while the CNN–Transformer fusion model in [vianto_2025] $51.33\%$. Using handcrafted fNIRS features, the ensemble classifier in [khan_aziz_2025] the authors attains $53.66\%$, while the video transformer in [nguyen_yang_2024] reports $55.00\%$. The strongest performance is presented by a multimodal transformer in [gkikas_rojas_painformer_2025], which achieves $55.69\%$ combining facial video and fNIRS.
The proposed method reaches $50.00\%$ accuracy on the test set with the Wave+PSD stack fusion. Even though it relies only on brain activity signals, it outperforms the SVM-based fNIRS baselines. It also remains competitive with several deep learning methods that include behavioral information, which supports the potential of unified tokenization and representation fusion for fNIRS-based pain recognition.
@@ VI Conclusion
In this study, we present a lightweight transformer for fNIRS-based pain recognition that combines unified tokenization with latent segmentation, allowing multiple signal representations to be processed by a single model. Experiments on the AI4Pain dataset showed that the proposed model maintained a stable performance across a range of computational settings. Fusing raw waveform and PSD representations demonstrated the feasibility of joint processing within a single backbone. These findings support the viability of efficient transformer designs for neurophysiological pain assessment in constrained settings [gkikas_tiny_2025, antonogiorgakis_britzolakis_2019], including clinical and interactive environments [10807220] where social robots could use real-time estimates of pain and cognitive states to adapt their behavior [arzate_when_2025, arzate_hri_2025, arzate_3dmm_2022, fang_moleron_2024, kruger_oshima_2026, hessels_fang_2026, vazquez_cruz_gkikas_2026].